OpenStudy (javaz):

The function ƒ(x) = −x2 − 9 is not one-to-one. Find a portion of the domain where the function is one-to-one and find an inverse function.

OpenStudy (javaz):

yes

OpenStudy (liziekay):

Do you know how to find an inverse function?

OpenStudy (javaz):

no

OpenStudy (will.h):

you may find the inverse function by switching the places of x and y and then solve for y like the following \[x = -y^2 - 9\] solve for y \[\sqrt{-x + 9} = f(^-1)\]

OpenStudy (will.h):

make sense so far?

OpenStudy (javaz):

where does the y go

OpenStudy (will.h):

y is the inverse but inverse symbol should be \[f^-1\]